An ac generator has emf epsilon = epsilon_m sin(om
Answer #1

a) emf is maximum when sin(wdt - pi/4) = 1

318t - pi/4 = pi/2

t = 0.75 pi/318 = 0.00741 s

b)current is

maximum when sin(wdt - 3pi/4) = 1

318t - 3pi/4 = pi/2

t = 1.75 pi/318 = 0.0173 s

c) it is inductor since current is lagging

d) XL =31.5/0.732 =43.033

L = 43.033/wd = 43.0327/318 = 0.135 H